OpenStudy (anonymous):

1,2,4,7,11 only the rule, and not in n's, i need it in y's and x's

OpenStudy (anonymous):

stupid teacher... asking for the wrong notation lol

OpenStudy (anonymous):

You know that all number ordering is in \[n\] right?

myininaya (myininaya):

why is y and x special?

OpenStudy (anonymous):

i cant figure out the rule here...

OpenStudy (anonymous):

well, thats how my teacher wants it, lawl

OpenStudy (anonymous):

tell her shes retarded

OpenStudy (anonymous):

its a he, lol

OpenStudy (anonymous):

lol i didnt see it at first either arman, myininaya got it though.

OpenStudy (anonymous):

oh nevermind

OpenStudy (anonymous):

how do you know minaya got it??

OpenStudy (anonymous):

n=(2n) +1

OpenStudy (anonymous):

that doesnt apply to the first term though

myininaya (myininaya):

i don't think he wants my recursive pattern i think he is rejecting it

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

mines wrong.. :P

OpenStudy (anonymous):

its a recursive pattern for sure, the rest of this guys questions were

OpenStudy (anonymous):

the last time he posted the question Myininaya got it, but i guess the teacher doesnt want that type of notation.

myininaya (myininaya):

i said \[f_n=f_{n-1}+n, where f_1=2\]

OpenStudy (anonymous):

maybe we could find a closed formula? one that isnt recursive?

myininaya (myininaya):

for n>=2

OpenStudy (anonymous):

thats not correct though, is it?

OpenStudy (anonymous):

fn=fn-1+n, where f1=2 is that what that says?

OpenStudy (anonymous):

oh damn it is, and its exatly what i was thinking too lol stupid fibancci

myininaya (myininaya):

\[f_2=f_{2-1}+2=f_1+2=2+2=4\] \[f_3=f_{3-1}+3=f_2+3=4+3=7\]

OpenStudy (anonymous):

tell your teacher to shutup and accept this as an answer

OpenStudy (anonymous):

i didnt know algebra homework was this complicated

myininaya (myininaya):

\[f_4=f_{4-1}+4=f_3+4=7+4=11\]

OpenStudy (anonymous):

it isint, ur teacher is just stupid...

OpenStudy (anonymous):

ignorant, asinine, naive, insolent, retarded, foolish, incapable, under-sophisticated, under-qualified, etc

OpenStudy (anonymous):

well the one I answered before the 2 4 6 8 16 one I got wrong canw e go back and work that one out

OpenStudy (anonymous):

i looked again and its actually bad

OpenStudy (anonymous):

the 2 4 6 8 16 one, is wrong?

OpenStudy (anonymous):

yeah sorry.. it works for the first 4 terms but then goes haywire with the 5th

OpenStudy (anonymous):

ugggh, why did i join honors algebra, i shouldve just done geometry, dumb summer homework

OpenStudy (anonymous):

i think this will work as a closed formula: the nth term is: \[\frac{1}{2}(n(n-1))+1\]

OpenStudy (anonymous):

how about: \[y _{n} = y _{n-1}+x\]

OpenStudy (anonymous):

its in xs and ys and its the same as the series formula

myininaya (myininaya):

lol does x=n?

OpenStudy (anonymous):

shutup...

OpenStudy (anonymous):

technically ANYTHING could equal n :p

myininaya (myininaya):

so mean

OpenStudy (anonymous):

now help me fix the other one!

OpenStudy (anonymous):

nuh uh! youre meaner :(

myininaya (myininaya):

no you

OpenStudy (anonymous):

no you

myininaya (myininaya):

:)

OpenStudy (anonymous):

okay, can you help me with the 2 4 6 8 16 one, cause this one is fine

OpenStudy (anonymous):

lol this is hilarious

OpenStudy (anonymous):

sigh your both as mean as each other now help me with the other one its driving me bonkers

myininaya (myininaya):

ok i'm going to bed you guys figure it out besides you are suppose to be more evolved in mathematics than us females have fun

OpenStudy (anonymous):

<.<

OpenStudy (anonymous):

myininaya=arman1002

OpenStudy (anonymous):

holy pellet... myininaya is a female!?!?!?!?!!?!?!?!?!?!?!!?!?!

OpenStudy (anonymous):

lmao so? Maybe I'M a female

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

omg... dude im being straight up serious... what the F

myininaya (myininaya):

lol

OpenStudy (anonymous):

are you a math teacher??

myininaya (myininaya):

yes

OpenStudy (anonymous):

professor?

OpenStudy (anonymous):

myininaya is awesome lets leave it at that

OpenStudy (anonymous):

omigawd... what grade you teach? calc?

myininaya (myininaya):

i teach at a small college

OpenStudy (anonymous):

aw dangit

OpenStudy (anonymous):

BAD ARMAN HELP WAYLANDER WITH 2,4,6,8,16!!

OpenStudy (anonymous):

anyways, waylanders formula works as long as 1 is the first term

OpenStudy (anonymous):

dude... 2,4,6,8,16 isint the sequence... or is it a new one we have to figure out?

OpenStudy (anonymous):

its one that has been 'answered' by me and its wrong

OpenStudy (anonymous):

uhh... 2,4,6,8,16... hmph, difficult one

OpenStudy (anonymous):

doesnt make sense to me...

OpenStudy (anonymous):

yeah i dont see it at all.

OpenStudy (anonymous):

unless its some kind of exponential

OpenStudy (anonymous):

it appears linear for the first 4 terms but then the 16 crops up and it is totally not linear

OpenStudy (anonymous):

it could be one of those gay ones lol